# Revolutionizing Future Connectivity: A Contemporary Survey on AI-empowered Satellite-based Non-Terrestrial Networks in 6G

### v3

@@ III-E1 Federated Learning (FL)
FL is a distributed learning technique in which multiple network data owners collaboratively build and train a global DL model, all while ensuring data isolation and privacy, as outlined in [101]. In the FL paradigm, individual data owners are initially provided with a base model by a centralized server, which they then train using their own respective data. Subsequently, these locally trained models are shared back with the central server, allowing it to update and maintain a global model. This iterative process continues until the global model reaches convergence. Consequently, FL enables the development of a globally trained model through distributed efforts, all the while safeguarding the privacy of the data owners.
@@ III-E2 Decentralized Learning (DcL)
DcL involves computing nodes conducting local training on their individual DL/ML models and then sharing these models with neighboring computing nodes at each iteration. Global convergence is achieved when all the local models have converged. Notably, this approach ensures that no actual data is exchanged between the computing nodes, but the local models are shared among neighboring nodes. One notable advantage of DcL is the absence of the need for a centralized server, which is a requirement in FL. An illustrative example of DcL can be found in the context of MARL, where agents collaboratively train their models in a distributed manner, as elaborated in [102].
@@ III-E3 Split Learning (SpL)
In SpL, instead of sharing model parameters, training occurs across various computing nodes, as described in [103]. Each computing node has the responsibility for training multiple layers of NNs within a DL model. Gradients for backpropagation are exchanged among these nodes to enhance training efficiency. Consequently, it can yield superior privacy performance when compared to FL, as indicated in [104]. Recent endeavors have been made to combine these two approaches, aiming to harness the advantages of both methodologies, as explored in [105].
@@ III-F Synergy between AI and NTN
Like many other fields, NTN is expected to be a major advancement in the realm of AI applications [106]. More precise and pragmatic analytical models with reduced overhead consumption, and efficient algorithms with a lower computational complexity are the primary catalysts for the deployment of AI-enabled NTN in next-generation wireless networks. In the preceding sections, we give a concise overview of NTN and AI to introduce these two crucial aspects of this article. Now, we motivate our readers by outlining the primary motivating forces behind combining AI and NTN for future wireless networks.
@@ III-F2 Tractable Solutions
The deployment of next-generation NTNs is more complex than any other previous-generation cellular network due to its multifaceted architecture. For instance, the integration of satellite networks introduces a significant number of additional parameters to consider for optimum network performance [107]. However, this can result in computationally intractable solutions for practical networks, even if the solutions are computationally tractable, they may be very inefficient. Resource management in TNTN networks is a prime example of this, as resource optimization in TNTN networks often turns into non-convex optimization problems, where only suboptimal or heuristic solutions can be obtained using numerical techniques [108, 109]. Fortunately, DL techniques can approximate complicated functions involving a large number of input variables with the help of NNs, as discussed in Section III-C. As a result, complicated network functionalities can be characterized with NNs and resource management issues can be solved in a tractable manner [110, 111].
@@ III-F6 Reduced Transmission Overhead
In some cases, traditional methods heavily rely on the exchange of information between various network participants, such as satellites and users. This might lead to a large overhead in communication channels, resulting in a decrease in the overall throughput of the network. AI can be used to reduce the control overhead of NTNs significantly. For example, to calculate the Doppler shift, the UEs must be provided with the latest ephemeris information of the satellites [112]. However, this would cause an immense overhead and a decline in the achievable data rate for the UEs. Alternative DL techniques can be employed to estimate the Doppler shift without requiring any ephemeris information from the satellites [113]. This leads to a significant decrease in transmission overhead over communication channels, resulting in superior network throughput.
@@ III-F7 Real-time Implementation
Network optimization and management decisions in NTNs usually require real-time implementation, usually in the order of milliseconds to tens of milliseconds. Consequently, complex algorithms cannot be used to obtain these real-time decisions. In most cases, the algorithms become either heuristic or offline. To have an online adaptable approach, AI techniques can be considered as a suitable option. For example, an online DRL-based approach as discussed in Section III-B can be used to obtain resource management decisions in real-time and ensure proper utilization of available resources in NTNs [114]. This is particularly valuable in latency-sensitive decision-making, such as scheduling, handover decisions, etc.
@@ IV Related Works
| year |
NTN challenges |
AI relevance |
Research thrusts |
6G perspective |
Current efforts |
Practical challenges |
[115] |
2021 |
[19] |
2021 |
[116] |
2021 |
[117] |
2022 |
[118] |
2022 |
[119] |
2022 |
[120] |
2023 |
[121] |
2020 |
[122] |
2020 |
[123] |
2020 |
[124] |
2021 |
[125] |
2022 |
[126] |
2023 |
[127] |
2019 |
Short |
[106] |
2019 |
Short |
Short |
[128] |
2020 |
Comprehensive, but only covers IoT applications |
Short and only covers IoT applications |
[119] |
2021 |
Short |
Complete |
[115] |
2021 |
Short uncategorized |
Does not cover AI for NTN |
[129] |
2021 |
Comprehensive, but does not cover the 6G perspective |
[107] |
2022 |
Short |
Short |
[130] |
2023 |
Short |
[131] |
2023 |
Comprehensive, but only covers RL |
Only covers RL |
[132] |
2023 |
Comprehensive, but does not cover the 6G perspective |
The possibility of potential integration of NTNs into 5G-Advanced [14] and future 6G networks to support various future high-demanding use cases has attracted significant attention from the research community in recent times. This emerging area of research has spurred numerous investigations to address the unique challenges and opportunities posed by NTN integration. [120] discusses the potential integration aspects for satellites, which is an integral part of NTNs, into future communication networks. [19] presents a summary of 3GPP efforts towards supporting NTNs in the 5G-Advanced networks. [119] presents the real system prototypes along with the general overview discussion on NTNs. [115] presents the challenges from the aspects of different communication layers to provide better insights for addressing these issues. In [13, 14], a concise discussion on various NTN components, use cases, technological enablers, and challenges for realizing NTN in 6G is presented. In [133], a detailed survey on the evolution of satellite networks towards the convergence with terrestrial networks from 3G to 6G along with the proposed architectures, use cases, and challenges is presented. In [118], future architectural options, use cases along the challenges associated with NTN-integrated 6G networks are explored. In [134], the necessary architectural evolution for integrating NTNs into 6G networks along with the challenges is discussed. [135] specifically focuses on the integrated Space-Air-Ground Integrated Network (SAGIN) in 6G while discussing the above topics in the context of NTNs. Another short magazine paper, [117] on NTN architectures, motivational use cases in 6G, necessary 5G NR modifications and future research directions is also in the literature. In [116], a detailed discussion on architectural options for integrating NTN into future 6G and the challenges associated with it is presented.
Likewise, AI has been acting as a driving force for various applications in wireless environments, especially in the last couple of decades; many surveys have been published on these topics recently [136, 137, 138, 139]. To facilitate the potential of AI in the 5G-Advanced and 6G environments several research articles and surveys are in the literature [140, 141, 18, 142]. In [121, 123, 126], some short surveys on the role of AI enabling 6G networks focusing on the vision, research opportunities, and challenges are presented. [122] discusses the explainability of AI to address various 6G challenges. In [124, 125], comprehensive surveys on vision, enabling technologies, and applications for AI on 6G are presented. Some relevant surveys are also published focusing on different aspects of AI-enabled 6G like pervasive network intelligence [143], green communications [144], privacy [145, 146], network access and routing [147]. As NTNs are expected to be integrated into the existing terrestrial environment for the development of 6G networks, it is clear that AI is expected to play a crucial role in this process. To unleash the full potential of AI to enable NTN in 6G, we need to have a clear understanding of the potential issues of NTN, we can gain insight into what AI tools can be useful down the road to resolve those issues.
There have been a few research articles capturing the key aspects of AI as an enabling technology for NTN in 6G in the recent past. In [119, 115], a short discussion on important applications of AI/ML in satellite-based NTN communication for 6G is provided along with the general discussion on NTN. In [107], several potential AI approaches for sustainable integrated Terrestrial and Non-Terrestrial Networks (TNTN) with a focus on maritime networking are discussed in a concise manner. In [127], a brief discussion of ML approaches to tackle different potential problems associated with integrated TNTNs is presented. In [106], it provides a short discussion on ML approaches for a limited number of issues related to next-generation mega-satellite networks. In [130], a compact discussion on different ML and DL techniques at various layers of the Open Systems Interconnection (OSI) model for NTN integration into existing 5G infrastructures is presented. Even though the above-mentioned works attempt to capture the role of AI in future 6G networks for enabling integrated TNTN environments, they are generally brief and do not provide a comprehensive overview of works in this particular domain. In [128], the potential role of AI techniques in the provision of NTN-based Intelligent Internet of Things (IoT) services is discussed; they do not focus on cellular environments for future integrated TNTN 6G networks. In [129], reviews of potential AI approaches for both broadcasting and communication satellites are provided. However, they do not focus on the issues related to NTN-integrated 6G networks, rather only focus on general satellite communication. In [131], a comprehensive review of the control approaches like coverage, spectrum, interference, and mobility management required by NTN platforms that are solved using RL formulations is presented, but they do not focus on other AI approaches related to prediction and estimation. A very recent comprehensive survey paper on ML and DL applications on satellite communications is published [132]. However, they do not discuss the current research efforts from the integrated 6G perspective and the potential challenges of applying ML and DL techniques in this domain.
@@ V-B1 Channel Estimation
Therefore, ML-based methods are increasingly being adopted by the research community and vendors, as a promising alternative for channel prediction. This channel estimation can potentially be turned into an SL problem by considering channel features such as distance, time delay, received power, azimuth Angle of Arrival (AoA) and Departure (AoD), elevation angle, Root Mean Square (RMS) Delay Spread, and frequency as inputs and CSI as output labels. In [148], the reciprocity property of the downlink and uplink channels in Time Division Duplexing (TDD) systems is considered. So the downlink channel is estimated from uplink CSI using an LSTM-based DL model. In [149], CSI is estimated from historical CSI data using a CNN-LSTM model. However, as channel estimation is a near-real-time process, low-complexity NNs such as ESNs need to be explored for realistic implementations. In [150], a denoising CNN is used to reduce the LS channel estimation error. In [151], a CSI prediction scheme is presented without utilizing any ephemeris information, rather only using past CSI feedback information leveraging GRUs with low prediction error. In [152], an auto-regressive integrated moving average of past CSIs is utilized to predict future CSIs where the order of the past ones is determined by an LSTM network. In [153], graph attention networks are used for cascaded channel estimation for Reflective Intelligent Surface (RIS) assisted satellite networks in IoT communications. In [154], future CSI information is predicted using k-Nearest Neighbour and MLP-based algorithms from past CSI and some correlated network metrics such as latency, terminal velocity, weather, and environment state, etc. which is later used to adapt the modulation and coding scheme for next timestamp. In [155], an RNN-based CSI compression technique is presented especially focusing on future SAGIN networks. An ANN is trained to estimate the fading at 40 GHz band exploiting the knowledge of its previous channel states in [156]. In [157], an LSTM-based CSI prediction framework is discussed to provide in future NTN-integrated 6G networks.
@@ V-B2 Doppler Shift Estimation
| SL | UL | RL |
RL Model |
DL Model |
[158] |
Averaged PSD |
FCNN |
[159] |
Modulation scheme, delay profiles, SNR |
CNN-LSTM |
[113] |
RSRP |
FCNN |
[160] |
CSI |
CNN |
@@
For LEO satellites, due to high mobility, This frequency offset is pretty significant (48 kHz with a center frequency of 2GHz [20]). Due to these frequency offsets, UEs tune to some different carrier frequencies from their originally assigned carrier frequencies. This may lead to ICI between multiple UEs as discussed in Section II-E.
There have been significant efforts to characterize the Doppler effect for LEO satellites since the launching of communication satellites. In [161], an equation for Doppler shift is derived for the simple case of LEO satellites with circular orbits in the equatorial plane and ground observing points on the equator. In [162], the Doppler shift is analytically derived assuming the trajectory of the satellite with respect to the earth by a great circle arc and the speed of the satellite as constant. In [112], the Doppler shift is characterized by considering a new orbit generator using different orbital parameters through a rigorous analysis. UEs with Global Navigation Satellite System (GNSS) can get the global positioning of satellites and estimate the amount of Doppler shift needed to be addressed for the next transmission slot [163]. However, this increases the cost and complexity which may not be feasible for ground UEs [31]. Additionally, The GNSS signals are weak, not ubiquitous, and susceptible to interference and spoofing. Recently, there have been also efforts to estimate the Doppler shift in LEO satellite systems using various other approaches, such as stochastic geometry [164], Maximum A Posteriori (MAP) [165], algebraic solutions [166], two-stage estimators consisting of time-varying Burg spectral analyzer and alpha-beta filter [167] etc. In [168], the Doppler shift is estimated using reference signals in more than one frequency position in Orthogonal Frequency Division Multiplexing (OFDM) carrier in a 5G integrated NTN system.
@@
In wireless communication systems, due to the mobility of the transceivers, the channel between the transceivers changes significantly resulting in received signal power variation and Doppler shift. So, intuitively, the CSI of this channel should contain information about the Doppler shift. This idea has been already explored in terrestrial networks to generate a model using ML[158, 113, 159]. The ground truth values or the labels are usually generated using the ephemeris information. Different channel characteristic variables like Rician K factor, azimuth AoA width, mean azimuth AoA and channel estimation errors are generated randomly, and averaged Power Spectral Density (PSD) is used as inputs with some preprocessing to a multi-layered FCNN to estimate the Doppler shift in [158]. In [113], RSRP values mapped from an ambiguity reducer are used to generate the weights for an MLP. In [159], different time and frequency domain signals with various modulation schemes, delay profiles, and Signal to Noise Ratio (SNR) have been used as inputs to a hybrid CNN-LSTM model to estimate the Doppler shifts. In NTN, the research in this domain is still at the early stage The estimated CSI is used as input to a CNN model to estimate the Doppler shift in [160]. In the future, other potentially efficient SL models can be also explored to generate the real-time accurate Doppler shift in an online manner. In table IV, we summarize the AI approaches for Doppler shift estimation in NTN. Even though the DL techniques are found to be useful in estimating Doppler shift using channel parameters, Doppler shift can be also estimated by analyzing the predictable trajectory of the satellites. Complexity analysis is required to justify the applicability of these DL architectures replacing the state of art methods in real systems.

### v4

@@ III-E1 Federated Learning (FL)
FL is a distributed learning technique in which multiple network data owners collaboratively build and train a global DL model, all while ensuring data isolation and privacy, as outlined in [fl1]. In the FL paradigm, individual data owners are initially provided with a base model by a centralized server, which they then train using their own respective data. Subsequently, these locally trained models are shared back with the central server, allowing it to update and maintain a global model. This iterative process continues until the global model reaches convergence. Consequently, FL enables the development of a globally trained model through distributed efforts, all the while safeguarding the privacy of the data owners.
@@ III-E2 Decentralized Learning (DcL)
DcL involves computing nodes conducting local training on their individual DL/ML models and then sharing these models with neighboring computing nodes at each iteration. Global convergence is achieved when all the local models have converged. Notably, this approach ensures that no actual data is exchanged between the computing nodes, but the local models are shared among neighboring nodes. One notable advantage of DcL is the absence of the need for a centralized server, which is a requirement in FL. An illustrative example of DcL can be found in the context of MARL, where agents collaboratively train their models in a distributed manner, as elaborated in [distl].
@@ III-E3 Split Learning (SpL)
In SpL, instead of sharing model parameters, training occurs across various computing nodes, as described in [split1]. Each computing node has the responsibility for training multiple layers of NNs within a DL model. Gradients for backpropagation are exchanged among these nodes to enhance training efficiency. Consequently, it can yield superior privacy performance when compared to FL, as indicated in [split-fl]. Recent endeavors have been made to combine these two approaches, aiming to harness the advantages of both methodologies, as explored in [split-fl-comb].
@@ III-F Synergy between AI and NTN
Like many other fields, NTN is expected to be a major advancement in the realm of AI applications [ai-satellite]. More precise and pragmatic analytical models with reduced overhead consumption, and efficient algorithms with a lower computational complexity are the primary catalysts for the deployment of AI-enabled NTN in next-generation wireless networks. In the preceding sections, we give a concise overview of NTN and AI to introduce these two crucial aspects of this article. Now, we motivate our readers by outlining the primary motivating forces behind combining AI and NTN for future wireless networks.
@@ III-F2 Tractable Solutions
The deployment of next-generation NTNs is more complex than any other previous-generation cellular network due to its multifaceted architecture. For instance, the integration of satellite networks introduces a significant number of additional parameters to consider for optimum network performance [IN_AI_NTN_1]. However, this can result in computationally intractable solutions for practical networks, even if the solutions are computationally tractable, they may be very inefficient. Resource management in TNTN networks is a prime example of this, as resource optimization in TNTN networks often turns into non-convex optimization problems, where only suboptimal or heuristic solutions can be obtained using numerical techniques [bh-it1, bh-gop]. Fortunately, DL techniques can approximate complicated functions involving a large number of input variables with the help of NNs, as discussed in Section III-C. As a result, complicated network functionalities can be characterized with NNs and resource management issues can be solved in a tractable manner [bh-dl, bh-dl2].
@@ III-F6 Reduced Transmission Overhead
In some cases, traditional methods heavily rely on the exchange of information between various network participants, such as satellites and users. This might lead to a large overhead in communication channels, resulting in a decrease in the overall throughput of the network. AI can be used to reduce the control overhead of NTNs significantly. For example, to calculate the Doppler shift, the UEs must be provided with the latest ephemeris information of the satellites [dop-acc]. However, this would cause an immense overhead and a decline in the achievable data rate for the UEs. Alternative DL techniques can be employed to estimate the Doppler shift without requiring any ephemeris information from the satellites [dop-ml-rsrp]. This leads to a significant decrease in transmission overhead over communication channels, resulting in superior network throughput.
@@ III-F7 Real-time Implementation
Network optimization and management decisions in NTNs usually require real-time implementation, usually in the order of milliseconds to tens of milliseconds. Consequently, complex algorithms cannot be used to obtain these real-time decisions. In most cases, the algorithms become either heuristic or offline. To have an online adaptable approach, AI techniques can be considered as a suitable option. For example, an online DRL-based approach as discussed in Section III-B can be used to obtain resource management decisions in real-time and ensure proper utilization of available resources in NTNs [ntn-ra-leo-rl1]. This is particularly valuable in latency-sensitive decision-making, such as scheduling, handover decisions, etc.
@@ IV Related Works
| year | NTN challenges | AI relevance | Research thrusts | 6G perspective | Current efforts | Practical challenges | |
| [COMST_AI_NTN_2] | 2021 | ||||||
| [5G-space] | 2021 | ||||||
| [NTN_8] | 2021 | ||||||
| [NTN_7] | 2022 | ||||||
| [NTN_4] | 2022 | ||||||
| [COMST_AI_NTN_1] | 2022 | ||||||
| [COMST_NTN_3] | 2023 | ||||||
| [ai-6g1] | 2020 | ||||||
| [ai-6g2] | 2020 | ||||||
| [ai-6g3] | 2020 | ||||||
| [ai-6g4] | 2021 | ||||||
| [ai-6g6] | 2022 | ||||||
| [ai-6g5] | 2023 | ||||||
| [IWC_AI_NTN_2] | 2019 | Short | |||||
| [ai-satellite] | 2019 | Short | Short | ||||
| [ai-iiot-ntn] | 2020 | Comprehensive, but only covers IoT applications | Short and only covers IoT applications | ||||
| [COMST_AI_NTN_1] | 2021 | Short | Complete | ||||
| [COMST_AI_NTN_2] | 2021 | Short uncategorized | Does not cover AI for NTN | ||||
| [AI_NTN_MAIN_1] | 2021 | Comprehensive, but does not cover the 6G perspective | |||||
| [IN_AI_NTN_1] | 2022 | Short | Short | ||||
| [ml-ntn] | 2023 | Short | |||||
| [rl-ntn] | 2023 | Comprehensive, but only covers RL | Only covers RL | ||||
| [ml-dl-sat] | 2023 | Comprehensive, but does not cover the 6G perspective | |||||
The possibility of potential integration of NTNs into 5G-Advanced [NTN_2] and future 6G networks to support various future high-demanding use cases has attracted significant attention from the research community in recent times. This emerging area of research has spurred numerous investigations to address the unique challenges and opportunities posed by NTN integration. [COMST_NTN_3] discusses the potential integration aspects for satellites, which is an integral part of NTNs, into future communication networks. [5G-space] presents a summary of 3GPP efforts towards supporting NTNs in the 5G-Advanced networks. [COMST_AI_NTN_1] presents the real system prototypes along with the general overview discussion on NTNs. [COMST_AI_NTN_2] presents the challenges from the aspects of different communication layers to provide better insights for addressing these issues. In [NTN_1, NTN_2], a concise discussion on various NTN components, use cases, technological enablers, and challenges for realizing NTN in 6G is presented. In [NTN_3], a detailed survey on the evolution of satellite networks towards the convergence with terrestrial networks from 3G to 6G along with the proposed architectures, use cases, and challenges is presented. In [NTN_4], future architectural options, use cases along the challenges associated with NTN-integrated 6G networks are explored. In [NTN_5], the necessary architectural evolution for integrating NTNs into 6G networks along with the challenges is discussed. [NTN_6] specifically focuses on the integrated Space-Air-Ground Integrated Network (SAGIN) in 6G while discussing the above topics in the context of NTNs. Another short magazine paper, [NTN_7] on NTN architectures, motivational use cases in 6G, necessary 5G NR modifications and future research directions is also in the literature. In [NTN_8], a detailed discussion on architectural options for integrating NTN into future 6G and the challenges associated with it is presented.
Likewise, AI has been acting as a driving force for various applications in wireless environments, especially in the last couple of decades; many surveys have been published on these topics recently [ai-w2, ai-w3, ai-w4, ai-w5]. To facilitate the potential of AI in the 5G-Advanced and 6G environments several research articles and surveys are in the literature [ai-w1, ai-w6, ai-wireless, fed-wireless]. In [ai-6g1, ai-6g3, ai-6g5], some short surveys on the role of AI enabling 6G networks focusing on the vision, research opportunities, and challenges are presented. [ai-6g2] discusses the explainability of AI to address various 6G challenges. In [ai-6g4, ai-6g6], comprehensive surveys on vision, enabling technologies, and applications for AI on 6G are presented. Some relevant surveys are also published focusing on different aspects of AI-enabled 6G like pervasive network intelligence [ni-6g], green communications [gc-6g], privacy [pr-ml-6g, pr-ml-ed-6g], network access and routing [nr-ml-6g]. As NTNs are expected to be integrated into the existing terrestrial environment for the development of 6G networks, it is clear that AI is expected to play a crucial role in this process. To unleash the full potential of AI to enable NTN in 6G, we need to have a clear understanding of the potential issues of NTN, we can gain insight into what AI tools can be useful down the road to resolve those issues.
There have been a few research articles capturing the key aspects of AI as an enabling technology for NTN in 6G in the recent past. In [COMST_AI_NTN_1, COMST_AI_NTN_2], a short discussion on important applications of AI/ML in satellite-based NTN communication for 6G is provided along with the general discussion on NTN. In [IN_AI_NTN_1], several potential AI approaches for sustainable integrated Terrestrial and Non-Terrestrial Networks (TNTN) with a focus on maritime networking are discussed in a concise manner. In [IWC_AI_NTN_2], a brief discussion of ML approaches to tackle different potential problems associated with integrated TNTNs is presented. In [ai-satellite], it provides a short discussion on ML approaches for a limited number of issues related to next-generation mega-satellite networks. In [ml-ntn], a compact discussion on different ML and DL techniques at various layers of the Open Systems Interconnection (OSI) model for NTN integration into existing 5G infrastructures is presented. Even though the above-mentioned works attempt to capture the role of AI in future 6G networks for enabling integrated TNTN environments, they are generally brief and do not provide a comprehensive overview of works in this particular domain. In [ai-iiot-ntn], the potential role of AI techniques in the provision of NTN-based Intelligent Internet of Things (IoT) services is discussed; they do not focus on cellular environments for future integrated TNTN 6G networks. In [AI_NTN_MAIN_1], reviews of potential AI approaches for both broadcasting and communication satellites are provided. However, they do not focus on the issues related to NTN-integrated 6G networks, rather only focus on general satellite communication. In [rl-ntn], a comprehensive review of the control approaches like coverage, spectrum, interference, and mobility management required by NTN platforms that are solved using RL formulations is presented, but they do not focus on other AI approaches related to prediction and estimation. A very recent comprehensive survey paper on ML and DL applications on satellite communications is published [ml-dl-sat]. However, they do not discuss the current research efforts from the integrated 6G perspective and the potential challenges of applying ML and DL techniques in this domain.
@@ V-B1 Channel Estimation
Therefore, ML-based methods are increasingly being adopted by the research community and vendors, as a promising alternative for channel prediction. This channel estimation can potentially be turned into an SL problem by considering channel features such as distance, time delay, received power, azimuth Angle of Arrival (AoA) and Departure (AoD), elevation angle, Root Mean Square (RMS) Delay Spread, and frequency as inputs and CSI as output labels. In [ntn-pl2], the reciprocity property of the downlink and uplink channels in Time Division Duplexing (TDD) systems is considered. So the downlink channel is estimated from uplink CSI using an LSTM-based DL model. In [ntn-pl1], CSI is estimated from historical CSI data using a CNN-LSTM model. However, as channel estimation is a near-real-time process, low-complexity NNs such as ESNs need to be explored for realistic implementations. In [csi-dncnn], a denoising CNN is used to reduce the LS channel estimation error. In [csi-gru], a CSI prediction scheme is presented without utilizing any ephemeris information, rather only using past CSI feedback information leveraging GRUs with low prediction error. In [csi-arima-lstm], an auto-regressive integrated moving average of past CSIs is utilized to predict future CSIs where the order of the past ones is determined by an LSTM network. In [csi-graph], graph attention networks are used for cascaded channel estimation for Reflective Intelligent Surface (RIS) assisted satellite networks in IoT communications. In [csi-ml], future CSI information is predicted using k-Nearest Neighbour and MLP-based algorithms from past CSI and some correlated network metrics such as latency, terminal velocity, weather, and environment state, etc. which is later used to adapt the modulation and coding scheme for next timestamp. In [channel-rnn], an RNN-based CSI compression technique is presented especially focusing on future SAGIN networks. An ANN is trained to estimate the fading at 40 GHz band exploiting the knowledge of its previous channel states in [csi-ann]. In [csi-lstm], an LSTM-based CSI prediction framework is discussed to provide in future NTN-integrated 6G networks.
@@ V-B2 Doppler Shift Estimation
| SL | UL | RL | RL Model | DL Model | |||
| [dop-ml1-ch] | Averaged PSD | FCNN | |||||
| [dop-ml2-ch] | Modulation scheme, delay profiles, SNR | CNN-LSTM | |||||
| [dop-ml-rsrp] | RSRP | FCNN | |||||
| [dop-ml-sat] | CSI | CNN | |||||
@@
For LEO satellites, due to high mobility, This frequency offset is pretty significant (48 kHz with a center frequency of 2GHz [3gpp1]). Due to these frequency offsets, UEs tune to some different carrier frequencies from their originally assigned carrier frequencies. This may lead to ICI between multiple UEs as discussed in Section II-E.
There have been significant efforts to characterize the Doppler effect for LEO satellites since the launching of communication satellites. In [dop-sim], an equation for Doppler shift is derived for the simple case of LEO satellites with circular orbits in the equatorial plane and ground observing points on the equator. In [dop-char], the Doppler shift is analytically derived assuming the trajectory of the satellite with respect to the earth by a great circle arc and the speed of the satellite as constant. In [dop-acc], the Doppler shift is characterized by considering a new orbit generator using different orbital parameters through a rigorous analysis. UEs with Global Navigation Satellite System (GNSS) can get the global positioning of satellites and estimate the amount of Doppler shift needed to be addressed for the next transmission slot [dop-gnss]. However, this increases the cost and complexity which may not be feasible for ground UEs [3gpp2]. Additionally, The GNSS signals are weak, not ubiquitous, and susceptible to interference and spoofing. Recently, there have been also efforts to estimate the Doppler shift in LEO satellite systems using various other approaches, such as stochastic geometry [dop-stoc], Maximum A Posteriori (MAP) [dop-map], algebraic solutions [dop-alg], two-stage estimators consisting of time-varying Burg spectral analyzer and alpha-beta filter [dop-blind] etc. In [dop-lin], the Doppler shift is estimated using reference signals in more than one frequency position in Orthogonal Frequency Division Multiplexing (OFDM) carrier in a 5G integrated NTN system.
@@
In wireless communication systems, due to the mobility of the transceivers, the channel between the transceivers changes significantly resulting in received signal power variation and Doppler shift. So, intuitively, the CSI of this channel should contain information about the Doppler shift. This idea has been already explored in terrestrial networks to generate a model using ML[dop-ml1-ch, dop-ml-rsrp, dop-ml2-ch]. The ground truth values or the labels are usually generated using the ephemeris information. Different channel characteristic variables like Rician K factor, azimuth AoA width, mean azimuth AoA and channel estimation errors are generated randomly, and averaged Power Spectral Density (PSD) is used as inputs with some preprocessing to a multi-layered FCNN to estimate the Doppler shift in [dop-ml1-ch]. In [dop-ml-rsrp], RSRP values mapped from an ambiguity reducer are used to generate the weights for an MLP. In [dop-ml2-ch], different time and frequency domain signals with various modulation schemes, delay profiles, and Signal to Noise Ratio (SNR) have been used as inputs to a hybrid CNN-LSTM model to estimate the Doppler shifts. In NTN, the research in this domain is still at the early stage The estimated CSI is used as input to a CNN model to estimate the Doppler shift in [dop-ml-sat]. In the future, other potentially efficient SL models can be also explored to generate the real-time accurate Doppler shift in an online manner. In table IV, we summarize the AI approaches for Doppler shift estimation in NTN. Even though the DL techniques are found to be useful in estimating Doppler shift using channel parameters, Doppler shift can be also estimated by analyzing the predictable trajectory of the satellites. Complexity analysis is required to justify the applicability of these DL architectures replacing the state of art methods in real systems.